The intensity of radiation emitted by sun has its maximum value at a wavelength of 510nm and that emitted by North Star has maximum value 350nm.If they behave like black bodies,ratio of surface temperature of sun to star is?

Grade:Select Grade

1 Answers

neptune1000
6 years ago
assuming radiation emmision corresponding to maximum intensity, Q=hv,v=frequency, Q=h*c/a, a=wavelength,c=speed of light in vaccum,h=planck’s constant. from stefan’s law,Q=o*T^4,o=stefan’s constant. Following these formulae take the ratio corresponding to the two wavelengths